# QuEST: Low-bit Diffusion Model Quantization via Efficient Selective Finetuning

### v3

@@ Abstract
The practical deployment of diffusion models still suffers from the high memory and time overhead. While quantization paves a way for compression and acceleration, existing methods unfortunately fail when the models are quantized to low-bits. In this paper, we empirically unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules. To alleviate the intensified low-bit quantization difficulty stemming from the distribution imbalance, we propose finetuning the quantized model to better adapt to the activation distribution. Building on this idea, we identify two critical types of quantized layers: those holding vital temporal information and those sensitive to reduced bit-width, and finetune them to mitigate performance degradation with efficiency. We empirically verify that our approach modifies the activation distribution and provides meaningful temporal information, facilitating easier and more accurate quantization. Our method is evaluated over three high-resolution image generation tasks and achieves state-of-the-art performance under various bit-width settings, as well as being the first method to generate readable images on full 4-bit (i.e. W4A4) Stable Diffusion. Code is available here.
@@ 1 Introduction
Diffusion models [6, 15, 32] have achieved great success in the field of image generation recently. Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency [4]. The first obstacle is the time-consuming denoising process which requires hundreds to thousands of inference time steps that slows down the generation speed drastically. The other one is the increasing model size driven by the need of better image fidelity and higher image resolutions. Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are carefully restricted.
Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously [9]. It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation. For example, 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically [23]. The benefits will increase consistently when the bit-widths are further lowered. Hence, low-bit quantization of diffusion models emerges as a viable approach for efficiency enhancement.
However, existing diffusion model quantization approaches fail under low-bit settings. These methods focus either on timestep-aware calibration data construction [19, 34] or quantization noise correction [12], targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs [31] and ViTs [2]. Such approaches overlook the intrinsic diffusion model mechanisms that are related with quantization, leading to inconsistencies between method deployment and model characteristics. Therefore, we intend to identify unique features in diffusion models that are associated with quantization.
In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the important and large values are rare and appear inconsistently. Conventional quantization methods tend to either approximate the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error. Diffusion models exhibit varying functions at distinct time steps [3], such as forming object outlines at early time steps and constructing details latter. Thus, preserving accurate temporal information is important during quantization. However, current methods struggle to represent extensive time steps with limited bit-width, leading to similar time embeddings across different time steps and results in error accumulation. Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks [7] and residual blocks [10]. Whereas previous works consider each model component as equally important and apply quantization uniformly, we reveal that certain layers are particularly sensitive to perturbations from quantization, while others are more resilient.
Table 1: Comparison with different frameworks.
@@
Addressing the revealed properties, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning). Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations in activations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters. Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization by reducing the significant values (large yet sparse), as shown in Figure 1(a). Following the idea of weight finetuning, we compare different bit-width models (Figure 1(b)) and identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property . Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(c). With less than 10% of the total parameters involved, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and requires only supervision from the full-precision model in a data-free manner. Our contributions are summarized as follows:
  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules. We also theoretically discuss why previous methods are expected to fail and justify the sufficiency of weight finetuning under low-bit settings.
@@
  • Addressing the identified properties, we introduce QuEST, a new finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency. An efficient activation quantizer is also proposed for time information estimation. Our method can be executed in a data-free and parameter-efficient manner.
@@
  • Experiments on three high-resolution image generation tasks over four models demonstrate the superiority of our method, achieving state-of-the-art performance under various bit-width settings. We are also the first method to generate reasonable images for full 4-bit Stable Diffusion.
The detailed related works on diffusion model inference and quantization are referred to appendix for space restriction. Table 1 is a compact summary that shows the unique capabilities of our method, where we are able to overcome the limitations of previous quantization methods. In comparison with EfficientDM [11], we include a brief discussion here: Technically, EfficientDM aims to finetune via low-rank weight representations [16]. However, it introduces additional parameters for each layer and requires extensive iterations of training on LoRA weights. Our method instead does not include additional parameters and only needs to finetune the model for limited iterations with the same batch size. We have also found that EfficientDM does not quantize the matrix multiplications in the attention mechanism, while we quantize the matrices to 8-bits. Moreover, EfficientDM always keeps the first three layers and the last layer as full 8-bit (weight and activation), as well as keeping skip-connections and some linear layers as full-precision during implementation. Our method instead only kept the first and last layer as 8-bit and quantize all other layers to the required bit-width.
Figure 1: Overview of our findings and method. (a) Illustration of property and a way to mitigate the quantization difficulty. (b) Demonstration of how property and property are identified via comparing different bit-width models. (c) Framework of the proposed method at time step $t$. To address the quantization difficulty and achieve more quantization-friendly activation distribution efficiently, we reduce significance by selectively and progressively updating partial model layers’ weights. $e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively. While $\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$ are selected layer weights to be optimized using Equation 2.3.2, $\mathbf{W}_{l^{{}^{\prime}}}$ and $\mathbf{W}_{l^{{}^{\prime\prime}}}$ are frozen during finetuning. More details are included in Section 2.3.2.
@@ 2.1 Preliminaries
This is denoted as the dequantization process. The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’). Equation 1 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship [22]. While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic mechanisms in quantized diffusion models. In the next section, we investigate three properties of diffusion models that relate to quantization performance and show why typical quantization methods tend to fail.
2.2 Quantization-aware Properties of Diffusion Models
Property : The activation outputs are likely to have a majority of values close to zero, with the other values being numerically large and important. Previous works [19, 34] paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself. In Figure 2, we cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization. We find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse, which we term as "significant values" (circled in green). Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7]. More visualizations can be found in Figure 6. This phenomenon poses difficulties in minimizing the clipping error and is unfriendly for effective quantization.
@@ 2.2 Quantization-aware Properties of Diffusion Models
Figure 2: Illustration of imbalanced activation distributions. In the activation distribution of the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization. Our method reduces the significant values to enable easier quantization for improved performance.
@@
Moreover, we observe that these significant values are important for generation performance preservation (see Figure 7). With the significant values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time. Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped significant values, generating corrupted images. This inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions, as depicted in Figure 1(a).
Table 2: Time embedding type comparison.
Bit-width Time Embedding FID $\downarrow$ sFID $\downarrow$ W8A8 PTQ Quantized 7.58 22.07 Full Precision 6.77 22.03 Ours 5.61 21.22 W4A8 PTQ Quantized 8.59 22.74 Full Precision 7.55 21.69 Ours 6.95 23.17
Property : Accurate temporal information is critical for quantization performance. Diffusion model predictions are highly correlated with time step information. Specifically, time steps are transformed and projected into time embeddings and then added onto the image features, guiding the model to denoise correctly. Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms [39]. Under W8A8 and W4A8, quantizing the time embeddings leads to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively. This indicates that inaccurate time embedding quantization can lead to large degradation in generation performance, yielding even larger corruptions under lower bit-widths. Due to the different focus diffusion models have on image generation at different time steps, we infer that inaccurate time embeddings can cause mismatched input image and model functionality, resulting in possible oscillations in the sequence of noise removal. Thus, preserving diffusion model performance under quantization requires accurate temporal information.

### v4

@@ Abstract
The practical deployment of diffusion models is still hindered by the high memory and computational overhead. Although quantization paves a way for model compression and acceleration, existing methods face challenges in achieving low-bit quantization efficiently. In this paper, we identify imbalanced activation distributions as a primary source of quantization difficulty, and propose to adjust these distributions through weight finetuning to be more quantization-friendly. We provide both theoretical and empirical evidence supporting finetuning as a practical and reliable solution. Building on this approach, we further distinguish two critical types of quantized layers: those responsible for retaining essential temporal information and those particularly sensitive to bit-width reduction. By selectively finetuning these layers under both local and global supervision, we mitigate performance degradation while enhancing quantization efficiency. Our method demonstrates its efficacy across three high-resolution image generation tasks, obtaining state-of-the-art performance across multiple bit-width settings. Code is available at https://github.com/hatchetProject/QuEST.
@@ 1 Introduction
Diffusion models [6, 13, 28] have recently achieved remarkable success in image generation. However, this success comes at the cost of two major obstacles that limit their efficiency [3]. The first obstacle is the denoising process which requires hundreds to thousands of inference time steps, slowing down the generation speed drastically. The other is the increasing model size, driven by demands for better image fidelity and higher image resolutions. Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are carefully restricted.
Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously [8], making it a natural solution for deploying diffusion models efficiently. It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation. For example, 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and memory reduction theoretically [22]. Hence, low-bit quantization of diffusion models emerges as a viable approach for efficiency enhancement. Unfortunately, existing diffusion model quantization methods that perform well at higher bit-widths face significant limitations in low-bit settings: some only adjust the quantization parameters and fail under low-bit conditions [18, 30, 10], while others succeed but require substantial computational resources comparable to training a diffusion model from scratch [21, 32]. In this work, we aim for efficient low-bit quantization, thereby circumventing the latter choice of resource-intensive training.
@@
Table 1: Comparison with different frameworks. Our method is both efficient and effective for low-bit diffusion model quantization, also achieving a reduced overall bit-width.
We first reveal the current challenge within diffusion models that impede the effectiveness of current efficient low-bit quantization methods [18, 31, 15]. As illustrated in Fig. 1(a): activation distributions tend to be imbalanced, with most values clustering near zero, while essential high-magnitude values are sparse and inconsistently distributed. Existing quantization methods [30, 18, 19] either approximate large and sparse values, inadequately estimating numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error. To overcome this challenge, we propose to adjust the activation distributions via weight finetuning, where its feasibility is justified both theoretically and empirically. Nevertheless, finetuning the entire diffusion model is a highly computationally-expensive and time-consuming process, requiring over 80GB memory and numerous hours [5, 21]. Thus, developing an efficient finetuning strategy tailored for diffusion model quantization is important.
To facilitate efficient quantization, we further identify two key properties of quantized diffusion models that unlock new opportunities: diffusion models exhibit varying functions at distinct time steps [2], therefore preserving accurate temporal information is important during quantization; and diffusion models possess complex network architectures, incorporating various types of modules. Whereas previous works consider each module as equally important and apply quantization uniformly, we reveal that certain modules are particularly sensitive to perturbations from quantization, while others are more resilient.
Figure 1: Overview of our observations and method. (a) Illustration of the challenge we identified in low-bit diffusion model quantization and a potential solution. We propose to ease the quantization difficulty by refining the activation distribution to make it more quantization-friendly. (b) Overview of property and property , whose importance are identified based on their impact on the model’s generation performance, making them suitable module candidates for efficient finetuning. (c) Framework of the proposed method. $\textbf{W}_{\text{TE}},\textbf{W}_{\text{A}},\textbf{W}_{\text{F}}$ are the weights of the time embeddings layers, attention-related layers, and other frozen layers, respectively. s is the quantization parameter. To train with efficiency, we adopt a selective and progressive finetuning strategy, incorporating temporal layer alignment (TLA) and critical module alignment (CMA). A global loss is also used for network-level guidance, improving the generated image quality.
Based on the above findings, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning). Confronting the revealed quantization challenge, we first theoretically justify that weight finetuning can enhance model robustness toward large activation perturbations in low-bit settings, thereby reducing quantization error. In contrast, previous methods have struggled to properly balance clipping error and rounding error. Then we empirically demonstrate that by finetuning the model weights, the activation distributions are modified to be more amenable to quantization. As shown in Fig. 1(a), the activation distribution is adjusted by reducing the amount of large, sparse values and enhancing the compactness of the value distribution.
@@
Following the idea of weight finetuning, we compare the effects of quantizing different modules of diffusion models (Fig. 1(b)) and identify two types of layers as primary culprits to performance degradation: time embedding layers exhibiting property and attention-related layers associated with property . Consequently, we selectively and progressively finetune the small subsets of identified layers in conjunction with all activation quantization parameters, as illustrated in Fig. 1(c). The learning objective is crafted to align the quantized model with its full-precision counterpart at both local and global levels. Involving less than 7% of the total parameters, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and can be conducted in a data-free manner. Our contributions are summarized as follows:
  • We identify the current challenge in low-bit diffusion model quantization that hinders effective low-bit quantization, and propose to adjust the activation distributions via weight finetuning for easier quantization. Both theoretical and empirical discussions are provided to justify the feasibility of this strategy under low-bit settings.
@@
  • We uncover and validate two properties in quantized diffusion models as the main factors for degraded performance. Motivated by the identified properties, we introduce QuEST, a parameter-efficient finetuning strategy that trains the diffusion model selectively and progressively, achieving low-bit quantization capability with time and memory efficiency.
@@
  • Experiments on three high-resolution image generation tasks over four models demonstrate the superiority of our method, achieving state-of-the-art performance under various bit-width settings.
2 Related Works
2.1 Diffusion Model Inference
Diffusion models [13, 28] generate samples via an iterative denoising process. During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained through a denoising process:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}),$ (1)
where $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output. This denoising process in a typical diffusion model requires tens to thousands of iterations, making efficient inference extremely challenging. Practically, diffusion models typically adopt a UNet architecture [29], incorporating an encoder and a decoder. Usually, encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNet structures in latent diffusion models, in alignment with the other works.
2.2 Diffusion Model Quantization
Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation. The researches for diffusion model quantization fall into three categories: Quantization-Aware Training (QAT) [16, 21, 20], Post-Training Quantization (PTQ) [19, 24, 34, 30, 18, 35], and Parameter-Efficient Fine-Tuning methods [9]. QAT methods [21] train all parameters from scratch, being effective for low-bit quantization but are extremely resource-intensive. PTQ methods [33, 10, 15, 36, 31] calculate the quantization parameters based on a small calibration set, offering better efficiency. However, PTQ methods often rely on complex designs and fail at lower bit-widths. To achieve low-bit compatibility with high efficiency, Parameter-Efficient Fine-Tuning methods were proposed. The representative work EfficientDM [9] trains a low-rank adapter (LoRA) [14] for each layer to reduce training costs, and successfully scales to W4A4.
Our proposed method also adopts a parameter-efficient finetuning strategy, and differs from EfficientDM in the following aspects: Firstly, EfficientDM introduces extra weight parameters, requiring substantial training iterations on the LoRA weights. Our method instead does not include additional parameters. Secondly, EfficientDM does not quantize the matrix multiplications in the attention mechanism, as well as certain linear layers. Our method quantizes all layers, and is more time-efficient. Tab. 1 summarizes the differences between our method with the other works.
@@ 3.1 Preliminaries
This is denoted as the dequantization process. The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’). Eq. 2 indicates that the quantization error is composed of two factors: the clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship [19]. While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic characteristics in quantized diffusion models. In the following sections, we first examine the current challenges in diffusion model quantization and outline our finetuning motivation, providing theoretical justification. We then identify two properties that enable efficient finetuning, forming the basis of our proposed method.
@@
Figure 2: Illustration of the imbalanced activation distributions. In the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization. Our method refines the activation distributions by eliminating the large and sparse values, enabling easier quantization.
3.2 The Challenge in Low-bit Quantization
Previous works [18, 15] primarily address the varying activation distributions across different time steps, facilitating diffusion model quantization at higher bit-widths. However, these methods experience failure in low-bit settings. To investigate the potential reason, we focus on the activation distribution itself and conduct a layer-wise analysis, revealing the following challenge in full-precision diffusion models that impedes effective quantization:
Challenge: Despite the majority of values being close to zero in the activation outputs, there exist numerically large and sparse values holding significant importance.
Fig. 2 provides a detailed analysis, where activation values are clustered into uniformly distributed bins. We find that in some layers, though the majority of values are close to zero, there exist values that are relatively large and diverse (circled in green). Take the bin plot on the left as an example, the original activation values (blue line) range from [-10, 34] but with most values between [-0.6, 1.7]. Visualizations for more layers can be found in Appendix 9. This phenomenon poses difficulties in minimizing the clipping error and is unfriendly for effective quantization.
@@ 3.2 The Challenge in Low-bit Quantization
Moreover, these large and sparse values are important for generation performance preservation. We find that when replacing the few tokens with maximum values by random noises, the generated images’ quality is critically degraded (as shown in Appendix 10). With these large values being important and the small values appearing frequently, neither of them is negligible and needs to be carefully quantized at the same time. Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped values, generating corrupted images. This inspires us to refine the activation distributions to attain more quantization-friendly distributions, as depicted in Fig. 1(a).
However, the activation distribution cannot be directly manipulated. To address this, we instead finetune the model weights under quantization constraints, producing a new yet similar full-precision model whose quantized counterpart maintains performance comparable to the original full-precision model. Our experiments interestingly reveal that the proposed finetuning strategy effectively eliminates the large and sparse values (Fig. 2), reducing quantization difficulty. We detail our approach in Sec. 3.3 and provide further theoretical analysis in Sec. 3.4.
3.3 Quantization via Efficient Selective Finetuning
In this section, we introduce QuEST, an efficient finetuning method for diffusion models that can significantly boost low-bit performance with less time and memory usage. We also present the two unique properties in quantized diffusion models, which serve as the foundation for the design of our method. Fig. 1(c) illustrates our approach.
3.3.1 Data-free Efficient Network-wise Training.
We first present the general training pipeline of our method. To alleviate the need for substantial training data, we construct the calibration set in a data-free manner. By feeding random Gaussian noises $x_{T}$ into the full-precision model and sampling over different time steps, we can obtain the calibration data needed for finetuning the quantized model. In practice, we only have to infer the full-precision model a few times to gather the needed number of calibration samples, totaling 128 or 256 samples per time step.
As depicted in Fig. 1(c), to overcome the quantization challenge efficiently, we update partial model weights ($\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$) that only account for a small subset of parameters related to the time step $t$. The remaining weight parameters $\mathbf{W}_{\text{F}}$ are kept frozen during optimization. We also fix the weight quantization parameters during training, reducing the amount of parameters that need to be optimized. For instance, in LDM-4 [28], no more than 7% of the parameters are adjusted. The choices for the weights to be finetuned will be discussed in the following sections.
The activation quantization parameters can be viewed as additional model parameters. Therefore, we further propose a network-wise training strategy. Different from quantization methods using layer-wise or block-wise reconstruction [18, 30] that bind quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors together with the partial weight parameters. Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, we update the required parameters at once. In this way, we significantly save the time and memory needed for quantization.
3.3.2 Temporal Layer Alignment
The inference process of diffusion models is highly dependent on the temporal information. Specifically, integer time steps are transformed into time embeddings through one or two linear layers, then added to the intermediate model features. Motivated by this observation, we make the following analysis that is consistent with previous works [15, 31]:
Property : Although time embeddings depend solely on time steps and are easily obtainable, precise temporal information is crucial for optimal quantization.
TE Setting (W8A8 Model) FID $\downarrow$ sFID $\downarrow$ TE Setting (W4A8 Model) FID $\downarrow$ sFID $\downarrow$ PTQ 7.58 22.07 PTQ 8.59 22.74 FP 6.77 22.03 FP 7.55 21.69 Ours 5.61 21.22 Ours 6.95 23.17
Table 2: Ablations on time embedding (TE) settings. Finetuning the TE layers with our method surpasses full-precision embeddings, while the latter outperforms standard quantized ones.
Tab. 2 provides an empirical justification, where we quantitatively show the performance drop when quantizing time embeddings to different bit-width. Under W8A8 and W4A8 bit-width settings, solely quantizing the time embeddings can lead to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively. We infer the reason is that inaccurate time embeddings can cause mismatched input and model functionality, resulting in possible oscillations in the sequence of noise removal. Previous works either propose to learn dynamic quantization parameters across different time steps through a simple network [31], or calibrate the time embedding layers and projection layers across all time steps [15]. We instead focus on finetuning the time embedding layers, adjusting fewer modules without introducing additional parameters. The results in Tab. 2 also suggest that our method can improve the quantization performance, even surpassing the full-precision baseline.
Concretely, in a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations. This implies that the time information operates independently from the primary network flow. Thus, we refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}$:
$\mathcal{L}_{\text{TLA}}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||O(t;\mathbf{w}_{l})-\tilde{O}(t;\mathbf{w}_{l},\mathbf{s}_{l})||^{2}],$ (4)
where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers. $O(t;\mathbf{w}_{l})$ is the intermediate activation of the full-precision model representing the ground truth, and $\tilde{O}(t;\mathbf{w}_{l},\mathbf{s}_{l})$ is the quantized activation. This objective function indicates that the chosen weight parameters are consistently updated across different time steps, so as to ensure robustness to diverse temporal inputs. Different from other methods [9, 15] that obtain different sets of quantization parameters for each time step, we only use a single set for varying time steps, improving time efficiency and memory storage.
3.3.3 Critical Module Alignment
While inaccurate time embedding quantization reduces performance under low-bit settings, it does not cause the complete generation failure observed in fully quantized models. Through careful layer-wise empirical study, we make the following observation:
Property : Not all activations respond equally to reduced bit-width, as different activations exhibit varying levels of sensitivity, with certain critical layers being especially sensitive to quantization.